three airports have a total of 213.6 million annual arrivals and departures. the second airport has 14.1 million more arrivals and departures than the third airport. the first airport has 1.5 million more arrivals and departures than the third airport. Find the number from each individual airport
first airport____
second airport_____
third airport_____

Answer by addingup(3677)

a+b+c = 213.6 (1)
b = c+14.1
a = c+1.5
Let's substitute for a and b in equation (1):
c+1.5+c+14.1+c = 213.6
3c+15.6 = 213.6
3c = 198
c = 66
b = 66+14.1 = 80.1
a = 66+1.5 = 67.5
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
66+80.1+67.5 = 213.6 we have the correct answer
J